# Statistical Taylor Expansion: A New and Path-Independent Method for Uncertainty Analysis 111

### v17

@@ 5.3 Distributional Hole
  • When the second derivative is negative, the resulted distribution is left-sided and Delta-like, such as when n=3,x=0.2n=3,x=-0.2, which is the mirror image of the distribution when $n=3,x=0.2$.
In each case, the transition from $x=0$ to $x=0.2$ is continuous for the resulting mean and deviation.
@@ 6 Matrix Calculations
Figure 14: Error deviations (z-axis) of adjugate matrix $\widetilde{\mathbf{M}}^{A}-\mathbf{M}^{A}$ as a function of matrix size (x-axis) and input noise precision (y-axis).
@@
Let $j_{1},j_{2}\dots j_{m}]_{n}$ denote a permutation of length $m$ from the vector $(1,2\dots n)$, and let $j_{1},j_{2}\dots j_{m}>_{n}$ for the ordered permutation of length $m$ [20]. Formula (6.1) and (6.2) gives the statistical Taylor expansion of the determinant of a square matrix $\mathbf{M}$ of size $n$, when the uncertainties of matrix elements are all independent of each other. In Formula (6.2), $\mathbf{M}_{<i_{1}\dots i_{m}>_{n},[j_{1}\dots j_{m}]_{n}}$ is a sub-matrix for $\mathbf{M}$, in which $i_{1}\dots i_{m}>_{n}$ contains the row indices, and $j_{1}\dots j_{m}]_{n}$ contains the column indices.
The square matrix whose element is $(-1)^{i+j}|m_{j,i}|$ is defined as the adjugate matrix [20] $\mathbf{M}^{A}$ to the original square matrix $\mathbf{M}$, in which $m_{j,i}$ is the element of $\mathbf{M}$ at row index $j$ and column index $i$. Each $m_{j,i}$ is an integer between $2^{8},+2^{8}]$ so that $\mathbf{M}^{A}$ can be calculated precisely using integer arithmetic. By adding Gaussian noise of deviation $\delta x$ to each $m_{j,i}$, $\widetilde{\mathbf{M}}$ is created. The value error of $\mathbf{M}^{A}$ is the difference between $\widetilde{\mathbf{M}^{A}}$ and $\mathbf{M}^{A}$, while the result deviation is calculated by Formula (6.2), so that $\widetilde{\mathbf{M}^{A}}-\mathbf{M}^{A}$ is a distribution test. Figure 14 shows that the error deviations of $\widetilde{\mathbf{M}^{A}}-\mathbf{M}^{A}$ are very close to $1$ except when the matrix size is $8$ and the input uncertainty is $\delta x=10^{-17}$. This abnormality is caused by floating-point rounding errors only at this point, when the element values exceed the significand range of 64-bit floating-point representation, to result in a error deviation of $3.2$. Figure 14 demonstrates a typical distribution test of error deviation versus input noise deviation $\delta x$ and another specific dimension, with ideal coverage for all region except when $\delta x\rightarrow 0$.
@@
Let $\mathbf{I}$ be the identity matrix for $\mathbf{M}$ [20]. Formula (6.3) show the relation of $\mathbf{M}^{A}$ and $\mathbf{M}$ which leads to the definition of inverse matrix $\mathbf{M}^{-1}$ in Formula (6.4) [20]. $\tilde{\mathbf{M}}\times\tilde{\mathbf{M}}^{A}-|\tilde{\mathbf{M}}|\mathbf{I}$ tests each element value so that it is a value test. Figure 15 shows the error deviation of a typical value test: the error deviation decreases linearly with increasing input uncertainty deviation $\delta x$, because the value should match unless at $\delta x\rightarrow 0$ when floating-point rounding errors dominate. The matrix size is the specific dimension in Figure 15.
Because an element of the original matrix $\mathbf{M}$ appears multiple times in Formula (6.4), its result variance is very complicated using Formula (2.3). For example, Formula (6.5) shows the simplest case for Formula (6.4): the first-order approximation of a 2x2 matrix. Contrary to traditional approach, statistical Taylor expansion uses Formula (6.4) for matrix inversion instead of Gaussian elimination [5], because logically, the result should be symmetric for all matrix elements, as what Formula (6.5) has demonstrated.
In Formula (6.4), $\mathbf{M}^{-1}$ is dominated by $1/|\mathbf{M}|$, suggesting that the precision of $\mathbf{M}^{-1}$ is largely determined by the precision of $|\mathbf{M}|$. Figure 16 shows that there is a strong linear correlation between conditional numbers and the corresponding determinant precision of matrices. As a reference, Figure 16 presents the Hilbert matrix [20] for each matrix size and shows that the Hilbert matrices also follow the linear relation between determinant precision and condition number. Thus, determinant precision can replace matrix condition number.
@@ 7 Mathematical Library Functions
Table 1: The result error deviation of selected basic functions using variance arithmetic when $\delta x>10^{-15}$. The error deviation for $(1\pm\delta x)^{c}$ can be improved to $0.998\pm 0.034$ if diverge regions are excluded. The error deviation for $\sin(x\pm\delta x)$ can be improved to $1.000\pm 0.010$ if pole regions are excluded.
Figure 17: Error deviation for $\sin(x\pm\delta x)$ as a function of $x$ and $\delta x$. The x-axis represents $x$ value between $\pi$ and $\pi$. The y-axis represents $\delta x$ value between $10^{-16}$ and $1$. The z-axis shows the corresponding error deviation.
Figure 18: Values and uncertainties of $\log(e^{x})-x$ and $e^{\log(x)}-x$ as functions of $x$, evaluated at $0.1$ increment. When $x$ is 2’s fractional such as $1/2$ or $1$, the result uncertainties are significantly smaller.
@@
Table 1 shows that Formula (2.15), (2.17), (2.20), and (2.22) provide almost perfect deviation for the result values which are calculated using corresponding math library functions by sampling from $x\pm\delta x$ in which $\delta x$ is Gaussian input noise deviation. Table 1 shows the results from Python, which differs slightly from those from either C++ or Java.
The result error deviation is similar to Figure 14 but with the specific dimension as $x$ in $e^{x\pm\delta x}$ and $\log(x\pm\delta x)$, or $c$ in $(1+\delta x)^{c}$. The coverage is only proper when $\delta x<10^{-15}$.
Figure 17 shows that the error deviation for $\sin(x+\delta x)$ is $1.000\pm 0.010$, except approaching 0 when $x=\pm\pi/2$ and $\delta x<10^{-8}$. Near a distributional pole, the input uncertainty is suppressed, resulting in zero error deviation. As expected, $\delta^{2}\sin(x)$ exhibits the same periodicity as $\sin(x)$. The numerical errors of the library functions $\sin(x)$ and $\cos(x)$ over a larger range of $x$ will be examined in greater detail in Section 10.
@@
  • Figure 19 shows that the error deviation for $(x^{p})^{1/p}-x$ is $0.56$, independent on neither $x$ nor $p$.
The reasons why $(x^{p})^{1/p}-x$ has larger value errors than $\log(e^{x})-x$, and $e^{\log(x)}-x$ has almost no value error, are not clear.
@@ 8.1 Moving-Window Linear Regression Algorithm
Figure 20: Result of fitting $\alpha+\beta\;Y$ to a time-series input $Y$ within a moving window of size 22+12*2+1. The x-axis indicates the time index. The y-axis on the left corresponds to the value of $Y$, $\alpha$, and $\beta$, while the y-axis on the right corresponds to the uncertainty of $\alpha$ and $\beta$. The uncertainty for $Y$ is fixed at $0.2$. In the legend, Unadjusted refers to results obtained by directly applying Formula (8.5) and (8.6) using variance arithmetic, whereas Adjusted refers to using Formula (8.5) and (8.6) for $\alpha$ and $\beta$ values but Formula (8.7) and (8.8) for their variances.
@@
Formula (8.1) and (8.2) provide the least-square line-fit of $Y=\alpha+\beta X$ between two set of data $Y_{j}$} and $X_{j}$}, where $j$ is an integer index identifying $(X,Y)$ pairs in the sets [5].
@@
In many applications, data set $Y_{j}$} denotes an input data stream where $j$ represents the time index or sequence index. $Y_{j}$} is thus referred to as a time-series input, with $j$ corresponding to $X_{j}$. A moving window algorithm [5] is applied within a small window centered on each $j$. For each calculation window, $X_{j}=-H,-H+1\dots H-1,H$} where $H$ is an integer constant specifying the half width of the window. This choice ensures $\sum_{j}X_{j}=0$, which simplifies Formula (8.1) and (8.2) into Formula (8.3) and (8.4), respectively [17].
@@
The values of $(\alpha_{j},\beta_{j})$ can be derived from the previous values $(\alpha_{j-1},\beta_{j-1})$, allowing Formula (8.3) and (8.4) to be reformulated into the progressive moving-window calculation given by Formula (8.5) and (8.6), respectively [17].
@@ 8.2 Variance Adjustment
When the time series contains uncertainty, directly applying Formula (8.5) and (8.6) results in a loss of precision since both formulas reuse each input multiple times, thereby accumulating the variance of that input with every reuse. To prevent this, $\alpha_{j}$ and $\beta_{j}$ should still be calculated progressively using Formula (8.6) and (8.5), respectively, while the variances should instead be computed using Formula (8.7) and (8.8), respectively. Formula (8.8) is not progressive because the progressive form of $\delta^{2}\beta_{j}$ is more expensive in computation than Formula (8.8).
@@
Figure 20 also presents the results of the moving window fitting of $\alpha+\beta\;Y$ versus the time index $j$. The fitted values of $\alpha$ and $\beta$ follow the expected behave, exhibiting a characteristic delay of $H$ in $j$. When (8.3) and (8.4) are applied to compute the uncertainties of $\alpha$ and $\beta$, both uncertainties increase exponentially with the time index $j$. In contrast, when Formula (8.3) and (8.4) are used exclusively for value calculation, while Formula (8.7) and (8.8) are applied for variance computation, the resulting uncertainties of $\alpha$ and $\beta$ are $\frac{\delta Y}{\sqrt{2H+1}}$, and $\frac{\delta Y}{\sqrt{\frac{H(H+1)(2H+1)}{3}}}$. Both are less than the input uncertainty $\delta Y$, due to the averaging effect of the moving window.
@@ 8.3 Unspecified Input Error
9 Regressive Generation of Sin and Cos
Figure 22: The uncertainties (y-axis) of $\sin(\pi j/2^{18})$ versus $j$ (x-axis) and $L$ using either Quart or regressive sine function (legend).
Formula (9.2) and Formula (9.3) calculate $\sin(\pi j/2^{L}),\cos(\pi j/2^{L}),j=0\dots 2^{L-2}$ repressively for regression order $L=0\dots 17$ starting from Formula (9.1). Formula (9.4) shows that such regression guarantees both $\sin(x)^{2}+\cos(x)^{2}=1$ and $\sin(2x)=2\sin(x)\cos(x)$, so that value errors will not accumulate when the regression order increases.
$\sin(0)=\cos(\frac{\pi}{2})=0;\;\;\;\sin(\frac{\pi}{2})=\cos(0)=1;$ (9.1)
Formula (9.2) is not suitable for computing $\sin(x)$ as $x\rightarrow 0$ because it suffers from behavior analogous to catastrophic cancellation, resulting in excessive uncertainty. In Figure 22, the Quart sine function exhibits uncertainties proportional to $\sin(\frac{\pi}{2^{L}})$, whereas the regression sine function shows increasing uncertainties as $x\rightarrow 0$. Unlike catastrophic cancellation in floating-point arithmetic, variance arithmetic uses very coarse precision to signal that the regression algorithm is unfit to compute $\sin(x\rightarrow 0)$.
@@ 10.1 DFT (Discrete Fourier Transformation)
For each signal sequence $h[k]$, where $k=0,1\dots N-1$, and $N$ is a natural number, the discrete Fourier transform (DFT) $H[n]$, for $n=0,1\dots N-1$, along with its inverse transformation, is defined by Formula (10.1) and (10.2), respectively [5]. As a convention, $k$ denotes time index for $h[k]$ as a waveform, while $n$ represents frequency index for $H[n]$ as a spectrum.
@@
To avoid the modeling errors inherent in DFT, only Formula (10.1) and (10.2) are used in this paper.
@@ 10.2 FFT (Fast Fourier Transformation)
When $N=2^{L}$, where $L$ is a natural number named FFT order, the generalized Danielson-Lanczos lemma can be applied to DFT to produce FFT [5].
  • For each output, each input is used only once, therefore no dependency problem arises when decomposing FFT into arithmetic operations such as Formula (2.10), (2.11), (2.12), and (2.13).
@@
  • The computational complexity is proportional to $L$, since increasing $L$ by 1 adds an additional step involving a sum of two multiplications.
  • Each step in the forward transformation doubles the variance, so the uncertainty deviation increases with the FFT order $L$ as $\sqrt{2}^{L}$. Because the reverse transformation divides the result by $2^{L}$, its uncertainty deviation decreases with $L$ as $\sqrt{1/2}^{L}$. Consequently, the uncertainty deviation for the roundtrip transformation is therefore $\sqrt{2}^{L}\times\sqrt{1/2}^{L}=1$.
  • The forward and reverse transformations are identical except for a sign difference, meaning that they are essentially the same algorithm, and any observed difference arises solely from the input data.
In normal usage, forward and reverse FFT transforms differ in their data prospective of time domain versus frequency domain:
  • The forward transformation converts a time-domain sine or cosine signal into a frequency-domain spectrum in which most values are zero, causing its uncertainties to grow more rapidly by cancellation.
  • In contrast, the reverse transformation spreads the precise frequency-domain spectrum (where most values are zero) back into a time-domain sine or cosine signal, causing its uncertainties to grow more slowly by spreading.

### v18

@@ 5.3 Distributional Pole
  • When the second derivative is negative, the resulting distribution is left-sided and Delta-like, such as when n=3,x=0.2n=3,x=-0.2, which is the mirror image of the distribution when $n=3,x=0.2$.
In each case, the transition from $x=0$ to $x=\pm 0.2$ is continuous for the resulting mean, deviation and error distribution.
@@ 6 Matrix Calculations
Figure 14: Error deviations (z-axis) of adjugate matrix $\widetilde{\mathbf{M}}^{A}-\mathbf{M}^{A}$ as a function of input noise precision (x-axis) and matrix size (y-axis).
@@
[…]
$=\sum_{m=1\dots n}\sum_{<i_{1}\dots i_{m}>_{n}}\sum_{[j_{1}\dots j_{m}]_{n}}\mathbf{M}_{<i_{1}\dots i_{m}>_{n},[j_{1}\dots j_{m}]_{n}}\prod_{i=1\dots m}^{i\in\{i_{1}\dots i_{m}\}}\tilde{x}_{i,p_{i}};$ (6.2)
@@
Let $j_{1},j_{2}\dots j_{m}]_{n}$ denote a permutation of length $m$ from the vector $(1,2\dots n)$, and let $j_{1},j_{2}\dots j_{m}>_{n}$ for the ordered permutation of length $m$ [20]. Let $\mathbf{M}$ be a square matrix of size $n$ with element $x_{i,j}\pm\delta x_{i,j},i,j=1,2,\dots n$ at row index $i$ and column index $j$. Fourmula (6.2) presents the Taylor expansion of the determinant for matrix $\mathbf{M}$ when the uncertainties of matrix elements are all independent of each other, whereas Formulas (6.3) and (6.4) gives the statistical Taylor expansion of the determinant $|\mathbf{M}|$ [20]. In Formula (6.4), $\mathbf{M}_{<i_{1}\dots i_{m}>_{n},[j_{1}\dots j_{m}]_{n}}$ is a sub-matrix for $\mathbf{M}$, in which $i_{1}\dots i_{m}>_{n}$ contains the row indices, and $j_{1}\dots j_{m}]_{n}$ contains the column indices [20].
[…]
@@
Let $\mathbf{I}$ be the identity matrix for $\mathbf{M}$ [20]. Formula (6.5) shows the relationship between $\mathbf{M}^{A}$ and $\mathbf{M}$ which leads to the definition of inverse matrix $\mathbf{M}^{-1}$ in Formula (6.6) [20]. $\tilde{\mathbf{M}}\times\tilde{\mathbf{M}}^{A}-|\tilde{\mathbf{M}}|\mathbf{I}$ tests each element value such that it is a value test. Figure 15 shows the error deviation of a typical value test: the error deviation decreases linearly with increasing input uncertainty deviation $\delta x$, because the values should match unless at $\delta x\rightarrow 0$ when floating-point rounding errors dominate. The matrix size is the algorithm specific dimension in Figure 15.
Because an element of the original matrix $\mathbf{M}$ appears multiple times in Formula (6.6), its variance, as obtained using Formula (2.3), is very complicated. For example, Formula (6.7) shows the simplest case for Formula (6.6): the first-order approximation of a 2x2 matrix. Contrary to conventional approach, statistical Taylor expansion uses Formula (6.6) for matrix inversion instead of Gaussian elimination [5] because logically, the result should be symmetric for all matrix elements, as demonstrated by Formula (6.7).
In Formula (6.6), $\mathbf{M}^{-1}$ is dominated by $1/|\mathbf{M}|$, suggesting that the precision of $\mathbf{M}^{-1}$ is largely determined by the precision of $|\mathbf{M}|$. Figure 16 shows that a strong linear correlation exists between conditional numbers [20] and the corresponding determinant precision of the matrices. As a reference, Figure 16 presents the Hilbert matrix [20] for each matrix size and shows that the Hilbert matrices also follow the linear relation between determinant precision and condition number. Thus, determinant precision can replace matrix condition number to quantify overall matrix stability [20], and Formula (6.7) further shows the stability of each matrix element.
@@ 7 Mathematical Library Functions
Table 1: The result error deviations of selected basic functions with Gaussian input noise $x\pm\delta x$ when $\delta x>10^{-15}$, obtained by variance arithmetic. The error deviation for $(1\pm\delta x)^{c}$ can be improved to $0.998\pm 0.034$ if diverging regions are excluded. The error deviation for $\sin(x\pm\delta x)$ can be improved to $1.000\pm 0.010$ if pole regions are excluded.
Figure 17: Error deviation for $\sin(x\pm\delta x)$ as a function of $x$ and $\delta x$. The x-axis represents $x$ values between $\pi$ and $\pi$. The y-axis represents $\delta x$ values between $10^{-16}$ and $1$. The z-axis shows the corresponding error deviations.
Figure 18: Values and uncertainties of $\log(e^{x})-x$ and $e^{\log(x)}-x$ as functions of $x$, evaluated in increment of $0.1$. When $x$ is 2’s fractional such as $1/2$ or $1$, the resulting uncertainties are significantly smaller because of floating-point representation.
@@
Table 1 shows that by sampling random input $\tilde{x}$ from Gaussian noise $x+\delta x$, Formulas (2.15), (2.17), (2.20), and (2.22) provide almost perfect characterization $f(x)\pm\delta f$ for the result distributions of f(x~)f($\tilde${x}) in which f(x~)f($\tilde${x}) is the corresponding math library functions. Table 1 shows the results from Python, which differs slightly from those from either C++ or Java.
The result error deviation is similar to that shown in Figure 14 but with the specific dimension as $x$ in $e^{x\pm\delta x}$ and $\log(x\pm\delta x)$, or $c$ in $(1+\delta x)^{c}$. The coverage is proper when $\delta x<10^{-15}$.
Figure 17 shows that the error deviation for $\sin(x+\delta x)$ is $1.000\pm 0.010$, except approaching 0 when $x=\pm\pi/2$ and $\delta x<10^{-8}$. Near a distributional pole, the input uncertainty is suppressed, resulting in zero error deviation. As expected, $\delta^{2}\sin(x)$ exhibits the same periodicity as $\sin(x)$. The numerical errors of the library functions $\sin(x)$ and $\cos(x)$ over a larger range of $x$ are examined in greater detail in Section 9.
@@
  • Figure 19 shows that the error deviation for $(x^{p})^{1/p}-x$ is $0.56$, dependent on neither $x$ nor $p$.
The reasons why $(x^{p})^{1/p}-x$ has larger value errors than $\log(e^{x})-x$, and $e^{\log(x)}-x$ has nearly no value error, are not clear. It could relate to floating-point representation which stores each value $x$ as $\log_{2}(x)$. Figure 5 also suggests that using linear approximation $e^{x\pm\delta x}$ and $\log(x\pm\delta x)$ have less error than $(x\pm\delta x)^{p}$.
@@ 8.1 Moving-Window Linear Regression Algorithm
Figure 20: Result of fitting $\alpha+\beta\;Y$ to a time-series input $Y$ within a moving window of size 22+12*2+1. The x-axis indicates the time index. The y-axis on the left corresponds to the value of $Y$, $\alpha$, and $\beta$, while the y-axis on the right corresponds to the uncertainty of $\alpha$ and $\beta$. The uncertainty for $Y$ is fixed at $0.2$. In the legend, Unadjusted refers to results obtained by directly applying Formulas (8.5) and (8.6) using variance arithmetic, whereas Adjusted refers to using Formulas (8.5) and (8.6) for $\alpha$ and $\beta$ values but Formulas (8.7) and (8.8) for their variances.
@@
Formulas (8.1) and (8.2) provide the least-square line-fit of $Y=\alpha+\beta X$ between two set of data $Y_{j}$} and $X_{j}$}, where $j$ is an integer index identifying $(X,Y)$ pairs in the sets [5].
@@
In many applications, data set $Y_{j}$} denotes an input data stream where $j$ represents the time index or sequence index. $Y_{j}$} is thus referred to as a time-series input, with $j$ corresponding to $X_{j}$. A moving window algorithm [5] is applied within a small window centered on each $j$. For each calculation window, $X_{j}=-H,-H+1\dots H-1,H$} where $H$ is an integer constant specifying the half width of the window. This choice ensures $\sum_{j}X_{j}=0$, which simplifies Formulas (8.1) and (8.2) into Formulas (8.3) and (8.4), respectively [17].
@@
The values of $(\alpha_{j},\beta_{j})$ can be derived from the previous values $(\alpha_{j-1},\beta_{j-1})$, allowing Formulas (8.3) and (8.4) to be reformulated into the progressive moving-window calculation given by Formulas (8.5) and (8.6), respectively [17].
@@ 8.2 Variance Adjustment
When the time series contains uncertainty, directly applying Formulas (8.5) and (8.6) results in a loss of precision since both formulas reuse each input multiple times, thereby accumulating the variance of that input with every reuse. To prevent this, $\alpha_{j}$ and $\beta_{j}$ should still be calculated progressively using Formulas (8.6) and (8.5), respectively, while the variances should instead be computed using Formulas (8.7) and (8.8), respectively. Formula (8.8) is not progressive because the progressive form of $\delta^{2}\beta_{j}$ is more expensive in computation than Formula (8.8).
@@
Figure 20 also presents the results of the moving window fitting of $\alpha+\beta\;Y$ versus the time index $j$. The fitted values of $\alpha$ and $\beta$ follow the expected behave, exhibiting a characteristic delay of $H$ in $j$. When (8.3) and (8.4) are applied to compute the uncertainties of $\alpha$ and $\beta$, both uncertainties increase exponentially with the time index $j$. In contrast, when Formulas (8.3) and (8.4) are used exclusively for value calculation, while Formulas (8.7) and (8.8) are applied for variance computation, the resulting uncertainties of $\alpha$ and $\beta$ are $\frac{\delta Y}{\sqrt{2H+1}}$, and $\frac{\delta Y}{\sqrt{\frac{H(H+1)(2H+1)}{3}}}$. Both are less than the input uncertainty $\delta Y$, because of the averaging effect of the moving window.
@@ 9.1 DFT (Discrete Fourier Transformation)
For each signal sequence $h[k]$, where $k=0,1\dots N-1$, and $N$ is a natural number, the discrete Fourier transform (DFT) $H[n]$, for $n=0,1\dots N-1$, along with its inverse transformation, is defined by Formulas (9.1) and (9.2), respectively [5]. As a convention, $k$ denotes the time index for $h[k]$ as a waveform, whereas $n$ represents the frequency index for $H[n]$ as a spectrum.
@@
To avoid the modeling errors inherent in DFT, only Formulas (9.1) and (9.2) are used in this study.
@@ 9.2 FFT (Fast Fourier Transformation)
When $N=2^{L}$, where $L$ is a natural number called the FFT order, the generalized Danielson-Lanczos lemma can be applied to DFT to produce FFT [5].
  • For each output, each input is used only once, therefore no dependency problem arises when decomposing FFT into arithmetic operations such as Formulas (2.10), (2.11), (2.12), and (2.13).
@@
  • The computational complexity is proportional to $L$, because increasing $L$ by 1 adds an additional step involving a sum of two multiplications.
  • Each step in the forward transformation doubles the variance; hence the uncertainty deviation increases with the FFT order $L$ as $\sqrt{2}^{L}$. Because the reverse transformation divides the result by $2^{L}$, its uncertainty deviation decreases with $L$ as $\sqrt{1/2}^{L}$. Consequently, the uncertainty deviation for the round-trip transformation is therefore $\sqrt{2}^{L}\times\sqrt{1/2}^{L}=1$.
  • The forward and reverse transformations are identical except for a sign difference, implying that they are essentially the same algorithm, and any observed difference arises solely from the input data.